>
Mathematics
List of top Mathematics Questions
Let $$ I_n = \int (\cos^n x + \sin^n x) \, dx,\quad \text{and} \quad \frac{n - 1}{n} I_{n - 2} = \frac{\sin x \cos x}{n} f(x) $$ Then, what is $ f(x) $?
AP EAPCET - 2022
AP EAPCET
Mathematics
Integration
If \( a, b, c, d \ne 0 \) and \( \int \frac{at + b}{ct + d} dt = t \), then find \( \int f(x) dx \) where \( f(x) = \frac{ax + b}{cx + d} \)
AP EAPCET - 2022
AP EAPCET
Mathematics
Integration
\([ \cdot ]\) represents the greatest integer function. If \( \int_{\sqrt{\frac{18}{3}}}^{[x]} dx = a + b\sqrt{2} + c\sqrt{3} \), then find the value of \( a + b + c \).
AP EAPCET - 2022
AP EAPCET
Mathematics
Integration
Let $ T>0 $ be a fixed number. If $ f: \mathbb{R} \to \mathbb{R} $ is a continuous function such that $ f(x + T) = f(x) $, then: $$ \text{If } I = \int_0^T f(x)\, dx,\quad \text{then } \int_0^{5T} f(2x)\, dx = ? $$
AP EAPCET - 2022
AP EAPCET
Mathematics
Integration
Evaluate: $$ \int \frac{e^{\tan^{-1}x}}{1+x^2} \left[\left(\sec^{-1}(\sqrt{1+x^2})\right)^2 + \cos^{-1} \left( \frac{1 - x^2}{1 + x^2} \right)\right] dx $$
AP EAPCET - 2022
AP EAPCET
Mathematics
Integration
Five-digit numbers are formed using digits 1,2,3,4,5 without repetition. Then the probability that the randomly chosen number is divisible by 4 is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
From a well shuffled pack of 52 cards, two cards are drawn at random. Then, the probability of both the cards being kings is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
A manager decides to distribute Rs. 20000 between two employees X and Y. He knows X deserves more than Y, but does not know how much more. So, he decides to arbitrarily break Rs. 20000 into two parts and gives X the bigger part. Then, the chance that X gets twice as much as Y or more is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
There are 4 hotels in a town. If 3 men check into the hotels in a day, what is the probability that each checks into a different hotel?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
In a set of 30 game cards, 17 are white and the rest are green. Out of the 30, 4 white and 5 green are marked “IMPORTANT”. If a card is chosen randomly from this set, the possibility of choosing a green card or an “IMPORTANT” card is:
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
During the winter months, in a certain village in Scotland, the probability of a day having severe fog is 0.6. The probability that in a given week there will be exactly two days with severe fog is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
In a Binomial distribution $B(n, p)$, if the mean and variance are 15 and 10 respectively, then the value of the parameter $n$ is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
If 6 is the mean of a Poisson distribution, then
$P(X \geq 3) =$
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
In a box, there are 8 red, 7 blue, and 6 green balls. One ball is picked randomly. The probability that it is neither red nor green is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Two brothers X and Y appeared for an exam. Let:
- $ A $: event that X passed
- $ B $: event that Y passed
Given: $$ P(A) = \frac{1}{7},\quad P(B) = \frac{2}{9} $$ Assuming independence, find the probability that both passed, i.e., $ P(A \cap B) $.
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Person A can solve 90% of the problems given in the book and Person B can solve 70%. Then the probability that atleast one of them will solve the problem selected at random from the book is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Two dice are thrown simultaneously. What is the probability that the product of the numbers is even?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Three students \( X, Y, Z \) appear for an exam. Their passing probabilities are: \[ P(X) = \frac{1}{5},\quad P(Y) = \frac{1}{4},\quad P(Z) = 1 - \frac{2}{3} = \frac{1}{3} \] What is the probability that at least two of them pass?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
A box contains 100 balls, numbered from 1 to 100. If 3 balls are selected one after the other at random with replacement from the box, then the probability that the sum of the three numbers on the balls selected from the box is an odd number, is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
A discrete random variable \( X \sim B(n, p) \) (i.e., binomial distribution). Given: \[ P(X = 2) = P(X = 3) \] Then what is the mean of \( X \)?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Which of the following is \textbf{not} a property of a Binomial distribution?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
From a pack of 52 cards, 3 cards are drawn at random. What is the probability that one is an ace, one is a queen, and one is a jack?
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
The discrete random variables
$X$
and
$Y$
are independent from one another and are defined as
$X \sim B(16, 0.25)$
and
$Y \sim P(2)$.
Then the sum of the variances of
$X$
and
$Y$
is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
For two events $A$ and $B$, a true statement among the following is
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
The maximum value of the variance of a Binomial distribution with parameters \( n \) and \( p \) is:
AP EAPCET - 2022
AP EAPCET
Mathematics
Probability
Prev
1
...
1007
1008
1009
1010
1011
...
1778
Next